# FAID: Fine-Grained AI-Generated Text Detection Using Multi-Task Auxiliary and Multi-Level Contrastive Learning

### v2

@@ C.1 Experimental Setup
Unless noted otherwise, we use a batch size of 64, the AdamW optimizer with a learning rate of 2e52e-5 and weight decay of 1e41e-4, 50 epochs, and 2000 warm-up steps. For the fuzzy kNN component, we set top-K = 20. We use temperature = 0.7.
@@ C.2 Computational Cost
All experiments run on a single NVIDIA A100 (40 GB); total wall-clock training time for FAID on FAIDSet is approximately 5 hours.
In our setup using FAISS Douze et al. (2024) with CPU inference, the average query latency is approximately 5ms for FAIDSet with 84,000 embeddings on a standard server (Intel Xeon CPU, 64GB RAM). When performing with Fuzzy KNN, it takes some minor time to process, so the total average time for inference is 10ms.
@@ C.3 Real-world Management of Vector Database
In the case of FAIDSet, our entire training and validation embedding store requires approximately 200MB, which is easily handled by standard server disks.
@@
C.4 Generalizability of FAID on unseen human–LLM type of collaboration
To further assess FAID’s robustness to unseen collaborative writing patterns between humans and LLM systems, we conducted an additional experiment focusing on hybrid authorship styles that were not present in the original training dataset.
We performed a manual quality control process to enhance the reliability of our dataset and ensure the robustness of the FAID model. A team of five annotators, all IT-major university students fluent in English and aged 18-25, participated in the annotation process. Each annotator was assigned approximately 80 samples, covering all collaborative writing styles (LLM-generated, LLM-continued,LLM-paraphrased, and LLM-polished), resulting in a total of 400 human-reviewed instances representing various collaboration modes.
During the manual revision stage, annotators followed two key quality control principles:
  1. 1.
    Ensuring logical and informational consistency between the output samples and original texts, where the output length was controlled to be between 70%–150% of the source length.
  2. 2.
    Improving linguistic quality through spelling correction, synonym replacement, and contextual word refinement to ensure natural fluency and stylistic coherence.
FAID was evaluated on this manually curated dataset without any further fine-tuning. Despite these samples representing an unseen distribution, the model achieved a strong overall accuracy of 84.8%, precision of 82.8%, and recall of 85.0%, across all collaboration categories. This suggests a strong generalization capability beyond the data distributions encountered during model development. It also highlights FAID’s sensitivity to fine-grained stylistic blending, where human revision only partially conceals the generative footprint of LLMs.
C.5 Generalizability of FAID on real-world scenarios.
To further assess the generalizability of FAID beyond the controlled benchmark setting, we conducted an additional user study simulating realistic academic and professional writing scenarios. The goal was to evaluate whether FAID can maintain detection accuracy when applied to authentic, unconstrained text produced through real interactions with LLMs.
A group of five volunteers with diverse academic backgrounds (same as in the Appendix C.4). Each participant was instructed to engage with four popular AI systems: ChatGPT, Gemini, DeepSeek, and Llama 3.1 on Together.AI to generate text resembling authentic student or professional writing tasks. The interactions were open-ended but guided by three types of prompts commonly encountered in real-world use cases: writing a paper summary (as an abstract) and writing a passage for their own graduation thesis. Each participant was asked to generate five outputs per model, yielding a total of 200 real-world text samples (5 participants $\times$ 4 models $\times$ 10 trials). To ensure authenticity and diversity, participants were encouraged to adjust prompts iteratively, include follow-up clarifications, and perform light editing, mimicking realistic human–LLM co-writing behavior.
FAID achieved strong performance in this real-world setting, with an overall accuracy of 88.5%, precision of 85.9%, and recall of 89.7%, indicating that FAID still exhibits strong generalizability in uncontrolled real-world scenarios with in-domain settings.
@@ Appendix D Analysis on Similarity across Models and Model Families
To assess the stylistic and semantic consistency of AI-generated texts, we conducted a comprehensive analysis from various perspectives, including N-gram distribution, text length patterns, and semantic embedding visualization. These analyses allow us to examine similarities both within the same model family and across different model families, leading to a robust understanding of AI "authorship" characteristics.
@@ D.1 Text Distribution between LLM Families
We first examined the distribution of text length, measured in both word and character counts across outputs from five LLMs: Llama-3.3-70B-Instruct-Turbo Dubey et al. (2024), GPT-4o-mini OpenAI (2024), Gemini 2.0, Gemini 2.0 Flash-Lite, and Gemini 1.5 Flash Team et al. (2024). Using 2000 arXiv prompt seeds, each model generated a corresponding output, and their length distributions were plotted.
From Figure 3, distinct patterns emerged between families. Gemini models consistently produced shorter and more compact outputs, while Llama and GPT models demonstrated more variance and a tendency toward longer completions. Despite using different versions (e.g., Gemini 2.0 vs. Gemini 1.5 Flash), the Gemini outputs remained closely grouped in terms of both word and character counts, suggesting a shared output generation strategy and stylistic consistency across the family. In contrast, Llama and GPT distributions showed greater separation and variability.
This reinforces the hypothesis that text length behavior is not only model-dependent but also family-coherent, with Gemini models forming a distinct cluster.
@@ D.3 Embedding Visualization and Semantic Cohesion
To delve into deeper semantic alignment, we visualized embeddings generated by an unsupervised SimCSE XLM-RoBERTa-base model of generated texts from two model families: Gemini family (texts generated by Gemini 2.0, Gemini 2.0 Flash-Lite, and Gemini 1.5 Flash), and GPT-4o/4o-mini using PCA to reduce the high-dimensional embeddings into a space.
As shown in Figure 4, Gemini model embeddings form tight, overlapping clusters, indicating a high degree of semantic cohesion among their outputs. This clustering remains consistent across both sample sizes of 2000 texts. In contrast, GPT-4o/4o-mini embeddings occupy a separate region of the space, with greater dispersion and less overlap with Gemini clusters.
This visualization confirms that the Gemini family not only shares stylistic features but also maintains semantic coherence, distinguishing it from models of other families at a conceptual level.
@@ D.4 Conclusion
(a) Word length distribution across five LLMs.
(b) Character length distribution across five LLMs.
@@
(a) 2D embedding space.
(b) 3D embedding space.
Figure 4: Visualizations showing clustering behavior of Gemini model family (Gemini 2.0, Gemini 2.0 Flash-Lite, Gemini 1.5 Flash) and GPT-4o/4o-mini using 2D and 3D embeddings with sample size of 2000 texts.
@@
(b) Top 3-grams of Gemini 2.0 Flash (500 samples).
(c) Top 3-grams of Gemini 1.5 Flash (500 samples).
@@ Appendix E Model Selection for Detector
To identify the best encoder for our classification task, we evaluated each candidate model on both known generators (FAIDSet testing data) and the new unseen-generator test set introduced in Section 3. Each transformer-based encoder was fine-tuned on FAIDSet training data and then used to predict labels on the two evaluation splits. Table 10 summarizes the accuracy, F1-macro, Mean Squared Error (MSE), and Mean Absolute Error (MAE) for each model under both conditions.
@@
We first evaluated three popular monolingual models: RoBERTa‑base, Flan‑T5‑base, and e5‑base‑v2 using the same training and evaluation splits. RoBERTa‑base Liu et al. (2019) and e5‑base‑v2 Wang et al. (2022) achieved the most balanced trade‑off between classification accuracy and regression error (MSE, MAE), while Flan‑T5‑base Chung et al. (2022) lagged slightly in F1‑macro. These results indicate that a stronger encoder backbone yields more robust performance, motivating the exploration of multilingual variants for further gains.
@@
Next, we experimented with XLM‑RoBERTa‑base Conneau et al. (2019) and Multilingual‑e5‑base Wang et al. (2024a). Both models leverage cross-lingual pretraining, which, in our scenario, appears to improve the representation of diverse linguistic patterns present in FAIDSet. In particular, XLM‑RoBERTa‑base delivered a substantial boost in all metrics, suggesting that its multilingual training enhances generalization even on primarily monolingual inputs.
@@
Finally, we integrated contrastive learning via SimCSE Gao et al. (2021) to further refine the sentence embeddings. We tested both supervised (trained on NLI data) and unsupervised (trained on Wikipedia corpus) SimCSE variants applied to RoBERTa‑base. The unsupervised version outperformed its supervised counterpart, confirming prior findings that unsupervised SimCSE often yields stronger semantic encoders for downstream tasks. Encouraged by this, we adopted the unsupervised SimCSE approach on XLM‑RoBERTa‑base Vamvas and Sennrich (2023), which achieved the highest accuracy and lowest error rates across all configurations.
@@ F.1 The Need to Use a Vector Database
We first used the trained model to classify the text and observed a significant performance decline from 90% when evaluating on unseen domains or text generated by novel models. This drop reflects the model’s limited ability to generalize beyond the distribution of its original training data. To address this, we propose integrating a vector database that stores dense embeddings of all collected examples, both training and incoming unlabeled data.
By indexing and retrieving semantically similar examples at inference time, the vector database serves as a flexible, scalable memory that bridges gaps between training and test distributions, enhancing the classifier’s resilience to domain shifts and unseen generators. Specifically:
@@ F.2 Clustering Algorithm Selection
To improve our detector’s robustness against unseen domains and generators, we evaluated four clustering strategies on the encoded vectors stored in our vector database. Each algorithm was tasked with grouping text samples into human‐written, AI‐generated, and human–LLM collaborative categories, using both known‐generator data (held out from training) and entirely unseen generator data. Performance was measured in terms of accuracy and F1-macro score.
We encoded each text sample using the penultimate layer of our pre-trained sequence-classification model, then applied clustering within the vector database to assign soft or hard cluster labels corresponding to the three text-origin classes. From the results in Table 11, we draw some key findings:
  • Traditional algorithms show reasonable performance on held‐out known generators but degrade notably on unseen generators.
  • Fuzzy C‐Means leverages membership degrees to handle overlapping distributions, improving both accuracy and F1-macro by about 4% over k‐Means, with smaller degradation on unseen data.
  • Fuzzy KNN combines local neighbor information with fuzzy membership, achieving the best overall performance: 95.2% accuracy and 95.1% F1-macro on known generators, and 93.3% accuracy and 93.3% F1-macro on unseen generators.
Given its superior ability to adapt to novel domains and generators through weighted neighbor voting and soft cluster assignments, we adopt Fuzzy k-Nearest Neighbors as the clustering component in our whole architecture.

### v3

@@ C.1 Experimental Setup
Unless noted otherwise, we use a batch size of 64, the AdamW optimizer with a learning rate of $2\times 10^{-5}$ and weight decay of $10^{-4}$, 50 epochs, and 2000 warm-up steps. For the fuzzy kNN component, we set top-K to 20, and we use a temperature of 0.7.
@@ C.2 Computational Cost
We run all experiments on a single NVIDIA A100 (40 GB). The total wall-clock training time for FAID on FAIDSet is approximately 5 hours.
In our setup using FAISS Douze et al. (2025) with CPU inference, the average query latency is approximately 5ms for FAIDSet with 83k embeddings on a standard server (Intel Xeon CPU, 64GB RAM). When performing with Fuzzy KNN, it takes some small period of time to process, and thus the total average time for inference is 10ms.
@@ C.3 Real-world Management of Vector Database
For FAIDSet, our entire training and validation embedding store requires approximately 200 MB, which is easily handled by standard server disks.
@@ Appendix D Analysis on Similarity across Models and Model Families
To assess the stylistic and semantic consistency of AI-generated texts, we conducted a comprehensive analysis across multiple perspectives, including N-gram distributions, text length patterns, and semantic embedding visualizations. This allowed us to study the similarities within and across model families, leading to a robust understanding of AI “authorship” characteristics.
@@ D.1 Text Distribution between LLM Families
We analyzed the distribution of text length, measured by both word and character counts, across outputs from five LLMs: Llama-3.3-70B-Instruct-Turbo Dubey et al. (2024), GPT-4o-mini OpenAI (2024), Gemini 2.0, Gemini 2.0 Flash-Lite, and Gemini 1.5 Flash Gemini Team (2024). Using 2,000 arXiv prompt seeds, each model generated a single output, and we plotted the resulting length distributions.
The results are shown in Figure 3, where we can observe that clear family-level patterns emerge. Gemini models consistently produce shorter, more compact outputs, whereas Llama and GPT models exhibit greater variance and a stronger tendency toward longer completions. Despite differences across versions (e.g., Gemini 2.0 vs. Gemini 1.5 Flash), Gemini outputs remain tightly clustered in both word and character counts, indicating a shared generation strategy and strong stylistic consistency within the family. In contrast, Llama and GPT distributions show greater dispersion and inter-model variability.
This reinforces the hypothesis that text length is not only model-dependent but also family-coherent, with Gemini models forming a distinct cluster.
@@ D.3 Embedding Visualization and Semantic Cohesion
To explore semantic alignment in detail, we visualized the embeddings generated by an unsupervised SimCSE XLM-RoBERTa-base model on texts from two model families, Gemini and GPT, using PCA to project the high-dimensional embeddings into a lower-dimensional space for analysis.
As shown in Figure 4, Gemini model embeddings form tight, overlapping clusters, indicating a high degree of semantic cohesion and internal consistency among their outputs. This clustering behavior remains stable across both sample sizes of 2,000 texts, suggesting that the observed patterns are not driven by sampling artifacts. In contrast, GPT-4o/4o-mini embeddings occupy a distinct region of the embedding space, exhibiting greater dispersion and noticeably less overlap with the Gemini clusters.
Overall, this visualization confirms that the Gemini family not only shares stylistic features, but also demonstrates strong semantic coherence, effectively distinguishing it from models belonging to other families at a deeper conceptual level.
@@ D.4 Conclusion
(a) word length distribution across five LLMs
(b) character length distribution across five LLMs
@@
(a) 2D embedding space
(b) 3D embedding space
Figure 4: Visualizations showing clustering behavior of Gemini model family (Gemini 2.0, Gemini 2.0 Flash-Lite, Gemini 1.5 Flash) and GPT-4o/4o-mini using 2D and 3D embeddings with sample size of 2,000 texts.
@@
(b) top 3-grams of Gemini 2.0 Flash (500 samples)
(c) top 3-grams of Gemini 1.5 Flash (500 samples)
@@ Appendix E Model Selection for Detector
In order to identify the best encoder for our classification task, we evaluated each candidate model on both known generators (in the FAIDSet test data) and the new unseen-generators test set, which is introduced in Section 3.
Each transformer-based encoder was fine-tuned on FAIDSet training data and then used to predict labels on the two evaluation splits. Table 12 summarizes the accuracy, F1-macro, Mean Squared Error (MSE), and Mean Absolute Error (MAE) for each model under both conditions.
@@
We first evaluated three popular monolingual models: RoBERTa‑base, Flan‑T5‑base, and e5‑base‑v2 using the same training and evaluation splits. RoBERTa‑base Liu et al. (2019) and e5‑base‑v2 Wang et al. (2024a) achieved the most balanced trade‑off between classification accuracy and regression error (MSE, MAE), while Flan‑T5‑base Chung et al. (2022) lagged slightly in F1‑macro. These results indicate that a stronger encoder backbone yields more robust performance, motivating the exploration of multilingual variants for further gains.
@@
We next evaluated XLM-RoBERTa-base Conneau et al. (2020) and Multilingual-e5-base Wang et al. (2024b). Both models benefit from cross-lingual pretraining, which in our setting improves the representation of the diverse linguistic patterns present in FAIDSet. Notably, XLM-RoBERTa-base yields a substantial improvement across all evaluation metrics, indicating that its multilingual training enhances generalization even when applied to predominantly monolingual inputs.
@@
Finally, we incorporated contrastive learning via SimCSE Gao et al. (2021) to refine sentence embeddings. We evaluated supervised (trained on NLI data) and unsupervised (trained on the Wikipedia corpus) SimCSE variants applied to RoBERTa-base. The unsupervised variant outperformed its supervised counterpart, aligning with prior findings that unsupervised SimCSE produces stronger semantic encoders. Based on these results, we applied unsupervised SimCSE to XLM-RoBERTa-base Vamvas and Sennrich (2023), achieving the highest accuracy and lowest error rates.
@@ F.1 The Need to Use a Vector Database
We first applied the trained model to classify the input text and observed a substantial performance drop of 15–30% when evaluated on unseen data. This degradation underscores the model’s limited ability to generalize beyond its training distribution and its sensitivity to distributional shifts. To address this issue, we decided to integrate a vector database that stores dense embeddings of all examples, including both labeled instances and unlabeled data encountered during inference. By indexing and retrieving semantically similar examples during inference, the vector database serves as a flexible, scalable memory module that helps bridge the gap between the training and test distributions. This retrieval-based mechanism enhances the classifier’s robustness to domain shifts and unseen generators by grounding predictions in stylistically and semantically related examples.
Specifically:
@@ F.2 Clustering Algorithm Selection
To improve our detector’s robustness against unseen domains and generators, we evaluated four clustering strategies in our vector database. Each algorithm was tasked with grouping text samples into human‐written, AI‐generated, and human–LLM collaborative categories, using both known‐generator data (held out from training) and entirely unseen generator data. For evaluation, we used accuracy and F1-macro score.
We encoded each example using the penultimate layer of our classification model, then applied clustering within the vector database to assign soft or hard cluster labels corresponding to the three classes. The results are shown in Table 13, where we can see:
  • Traditional algorithms show reasonable performance on held‐out known generators, but degrade notably on unseen generators.
  • Fuzzy C‐Means leverages membership degrees to handle overlapping distributions, improving both measures by 4% over k‐Means, with smaller degradation on unseen data.
  • Fuzzy KNN combines local neighbor information with fuzzy membership, achieving the best overall performance.
Given its superior ability to adapt to novel domains and generators through weighted neighbor voting and soft cluster assignments, we adopt Fuzzy k-Nearest Neighbors as the clustering component in our overall architecture.